OpenStudy (anonymous):

Solve tanx = 2cosx*tanx

OpenStudy (anonymous):

Ok So first thing is to divide both sides by tanx \[ \frac{\tan x}{\tan x} =\frac { 2 \cos x * \tan x}{\tan x}\] \[1=2 \cos x\] \[\frac{1}{2}= \cos x \]

OpenStudy (anonymous):

Can you solve from here